# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v2

@@ 4 Approximation based on Hausdorff distance
In this section, we present our improved theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between the original and approximate neural networks. Misiakos et al. (2022) derived a bound for the error of approximating the tropical polynomials that represent a neural network. The motivation for geometrically bounding the error of these polynomials stems from Theorem 5. This theorem indicates that two polynomials with the same extended Newton polytopes are functionally equivalent. Consequently, it is expected that two tropical polynomials with approximately equal extended Newton polytopes will attain similar values.
@@
Definition 7 (Hausdorff distance)
Let $S,\tilde{S}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance $H(S,\tilde{S})$ of the two sets is defined as
In the case of polytopes, the following lemma (proof in Appendix A.2) reassures us that the suprema in the above expression are attained, and in fact by points in the vertex sets $V_{P},V_{\tilde{P}}$ of the polytopes.
Lemma 8
Due to the convexity and compactness of polytopes, we have that
In their work, Misiakos et al. (2022) used the discrete form of the Hausdorff distance, defined as the Hausdorff distance of the vertex sets of the two polytopes ($DH(P,\tilde{P}):=H(V_{P},V_{\tilde{P}})$), to obtain the following result.
Proposition 9 (Misiakos et al., 2022)
Let $p,\tilde{p}\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials with extended Newton polytopes $P=\mathrm{ENewt}(p)$ and $\tilde{P}=\mathrm{ENewt}(\tilde{p})$. Then,
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq DH(P,\tilde{P})$
@@
We extend the result of Misiakos et al. (2022). The proof is provided in the appendix.
Proposition 10
Let $p,\tilde{p}\in\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$ be two tropical polynomials with extended Newton polytopes $P=\mathrm{ENewt}(p)$ and $\tilde{P}=\mathrm{ENewt}(\tilde{p})$. Then,
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq H(P,\tilde{P})$
Since $V_{P}\subseteq P,V_{\tilde{P}}\subseteq\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$. The equality in the aforementioned inequalities is achieved only in special cases and thus the bound $H(P,\tilde{P})$ we provide is in general tighter than $DH(P,\tilde{P})$ from Misiakos et al. (2022). We demonstrate one example where the two formulations give equal distances and one where the standard Hausdorff distance is strictly smaller than the discrete one.
Example 2
Let $P$ be the zonotope formed by taking the generators $(1,0)$ and $(\cos 60^{\circ}=\frac{1}{2},\sin 60^{\circ}=\frac{\sqrt{3}}{2})$. Form zonotope $\tilde{P}$ by taking as a generator the mean of the two generators of $P$, which is the vector $(\frac{3}{4},\frac{2+\sqrt{3}}{4})$. Refer to Figure 4(a) for a visual representation.
The discrete Hausdorff distance between the two polytopes is given by:
$DH(P,\tilde{P})=\max\left\{\frac{\sqrt{3}}{2},\frac{1}{2}\right\}=\frac{\sqrt{3}}{2}.$
In this case, the Hausdorff distance is equal to the discrete one:
$H(P,\tilde{P})=\max\left\{\frac{\sqrt{3}}{2},\sin 30^{\circ}\cdot|\text{generator of }\tilde{P}|\right\}$
$=\max\left\{\frac{\sqrt{3}}{2},\frac{1}{2}\cdot\cos 30^{\circ}=\frac{\sqrt{3}}{4}\right\}=\frac{\sqrt{3}}{2}.$
Example 3
Let zonotope $P$ be formed by taking the generators $(1,0)$ and $(\cos 60^{\circ}=\frac{1}{2},\sin 60^{\circ}=\frac{\sqrt{3}}{2})$. Form zonotope $\tilde{P}$ by taking as a generator the sum of the two generators, which results in the vector $(\frac{3}{2},\frac{2+\sqrt{3}}{2})$. Refer to Figure 4(b) for a visual representation.
The discrete Hausdorff distance between the two polytopes is calculated as:
$DH(P,\tilde{P})=\max\{0,1\}=1.$
Here, the Hausdorff distance is strictly smaller and equal to:
$H(P,\tilde{P})=\max\left\{0,\frac{1}{2}\right\}=\frac{1}{2}.$
(a) Example 2
(b) Example 3
Figure 4: Polytopes $P$, $\tilde{P}$ of examples 2, 3
Notice that if our goal were to approximate the polytope $P$ by the polytope $\tilde{P}$, then according to the discrete Hausdorff distance, taking the mean is the optimal strategy and gives distance $\sqrt{3}/2$ (example 2), better than $1$ (example 3). On the other hand, according to the Hausdorff distance in its standard form, taking the sum is the optimal strategy and gives distance $1/2$ (example 2), better than $\sqrt{3}/2$ (example 3). This implies that an algorithm based on the Hausdorff distance would imply a better approximation of $P$ in the above example.
We derive the following result, which differs from (Misiakos et al., 2022) by using the standard form of the Hausdorff distance and thus providing a tighter bound. For the proof we apply the triangle inequality and then repeatedly Proposition 10 to each output of the network. The full proofs of Proposition 10 and Theorem 11 are provided in Appendix A.3.
Theorem 11
Let $v$ and $\tilde{v}$ be the outputs of two neural networks as in Figure 3. Then, the following inequality holds:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
@@ 5 Compression Algorithm
In this section we present our algorithm TropNNC for neural network compression. Our method, based on Theorem 11, is a structured neural network compression method that reduces the number of neurons in the hidden layers and produces a compressed network that is a functional approximation of the original. We differentiate between single-output and multi-output networks and introduce novel approaches for both scenarios. The main achievement of our method is to preserve the neural network as a function, achieving performance comparable to state-of-the-art frameworks, all without relying on training data samples.
@@ 5.1 Single Output
First, we focus on single output networks, as the one depicted in Figure 3, with $m=1$. The network has weight matrices $\mathbf{A}\in\mathbb{R}^{n\times(d+1)}$, $\mathbf{C}\in\mathbb{R}^{1\times n}$, and a single output $v$ with corresponding positive and negative zonotopes $P$ and $Q$. We aim to approximate this network with a smaller one that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$. Our goal is to ensure that $\tilde{v}(\mathbf{x})\approx v(\mathbf{x})$ for all $\mathbf{x}\in B$. According to Theorem 11, it is sufficient to choose the new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$ such that $\widetilde{P}\approx P$ and $\widetilde{Q}\approx Q$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
In the original network with $n$ hidden neurons, each neuron contributes a generator $|c_{i}|(\mathbf{a}_{i}^{T},b_{i})$, which can be either positive or negative. In the approximating network with $K$ hidden neurons, each neuron contributes a generator $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$. We aim to pick generators $|\tilde{c_{i}}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ such that the generated positive and negative zonotopes $\widetilde{P}$ and $\widetilde{Q}$ approximate the positive and negative zonotopes $P$ and $Q$, respectively.
In essence, we have translated the neural network approximation problem into a zonotope approximation problem—specifically, the task of approximating a zonotope with another zonotope that has fewer generators. This problem is known in the literature as zonotope order reduction (Yang and Scott, 2018). In our case, the approximation must happen in terms of the Hausdorff distance.
5.1.1 Algorithm for single output network
Misiakos et al. (2022) proposed the algorithm ”Zonotope K-means” for compressing single-output neural networks. Their approach uses K-means clustering on the set of positive and negative generators, and replaces the generators of each formed cluster by a single representative, the center of the cluster (i.e. the mean of the generators of the cluster).
In contrast, our approach also uses K-means to cluster the generators, but instead of taking the representative to be the center of the cluster, we take the representative to be the sum of the generators of the cluster. We show later that this compression produces a tighter approximation of the neural network compared to (Misiakos et al., 2022). We demonstrate this with Example 4 and prove this formally with Proposition 12 and Corollary 13. Our Algorithm 1 is depicted below.
@@ 5.1.1 Algorithm for single output network
[…]
@@
5:  For each constructed neuron $i$ set the output weight to $1$ if the neuron corresponds to a representative of a positive cluster ($i\in C_{+}$), otherwise set it to -1.
Example 4
Suppose we have a positive zonotope consisting of 3 positive generators $g_{1},g_{2},g_{3}$ like the pink zonotope (hexagon) in Figure 5(a), and suppose we have $K/2=2$. The Zonotope K-means algorithm of Misiakos et al. (2022) will form 2 clusters of generators, the black cluster which consists of a single generator $g_{1}$, and the pink cluster which consists of the pink generators $g_{2},g_{3}$. Zonotope K-means will take as cluster representatives the centers of the clusters. In our example, this will be the black generator $\tilde{g}_{1}=g_{1}$ for the black cluster, and the blue generator $\tilde{g}_{2}=\frac{g_{2}+g_{3}}{2}$ for the pink cluster, as in Figure 5(b). The blue zonotope is the zonotope generated by the representatives of the clusters (black and blue generators $\tilde{g}_{1},\tilde{g}_{2}$ respectively), and it is the approximation that Zonotope K-means uses to approximate the pink zonotope.
In contrast, TropNNC will also form the same 2 clusters of generators, but instead of taking the center, it will take the sum of the generators of each cluster as the cluster representative. In our example, these will be the black generator $\tilde{g}_{1}=g_{1}$ and the blue generator $\tilde{g}_{2}=g_{2}+g_{3}$ of Figure 5(c), which form a different blue zonotope which is used as the approximation of the pink zonotope.
In this example, one can visually verify that according to the Hausdorff distance, the approximation of TropNNC is better than the approximation of Zonotope K-Means.
(a) Original Zonotope
(b) Zonotope K-means
@@
Figure 5: Example execution of Zonotope K-means and TropNNC
The result of Example 4 can be generalized. Indeed, the following bound and its corollary hold (proofs in Appendix A.4 and A.5).
Proposition 12
@@
$\frac{1}{\rho}\max_{\mathbf{x}\in B}|u(\mathbf{x})-\tilde{v}(\mathbf{x})|\leq\sum_{i\in I}\min\{||c_{i}(\mathbf{a}_{i}^{T},b_{i})||,\delta_{max}\},$
@@ 5.2 Multi-output
Now we consider the case of multi-output networks as in Figure 3, for general $m\in\mathbb{N}$. We consider a network with output $\mathbf{v}=(v_{1},\ldots,v_{m})$ and corresponding positive and negative zonotopes $P_{j}$ and $Q_{j}$ for each output node $v_{j}$. An interesting property of these polytopes is that they share the directions $(\mathbf{a}_{i}^{T},b_{i})$ of their generators. For instance, output $v_{1}$ might have a positive generator $|c_{1i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $P_{1}$, while output $v_{2}$ might have a negative generator $|c_{2i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $Q_{2}$. These generators are parallel to each other, with common direction $(\mathbf{a}_{i}^{T},b_{i})$.
As before, we aim to approximate this network with a smaller network that has fewer neurons in the hidden layer, say $K<n$, resulting in a new hidden layer $\tilde{\mathbf{f}}=(\tilde{f}_{1},\ldots,\tilde{f}_{K})$ with new weight matrices $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$. Our goal is to ensure that $\tilde{\mathbf{v}}(\mathbf{x})\approx\mathbf{v}(\mathbf{x})$ for all $\mathbf{x}\in B$. As we have seen, this can be translated to having $\widetilde{P}_{j}\approx P_{j}$ and $\widetilde{Q}_{j}\approx Q_{j}$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
Notice that for each new neuron $\tilde{f}_{i}$, its input weight-bias $(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$ defines a generator direction for every output $\tilde{v}_{j}$, where $j\in[m]$. For this generator direction, the magnitudes of the generators change across different outputs by adjusting the output weight $\tilde{c}_{ji}$ for each $j\in[m]$, resulting in parallel generators for each output: $|\tilde{c}_{1i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}),\ldots,|\tilde{c}_{mj}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})$. These generators can be either positive or negative according to the sign of $\tilde{c}_{ji}$.
Thus, we start with original positive and negative zonotopes, with parallel generators, and aim to approximate them with new positive and negative zonotopes, also with parallel generators. We refer to this complex approximation problem as simultaneous zonotope approximation.
Example 5
Suppose we have a neural network with a single hidden layer, as in Figure 3, with dimensions $d+1=n=m=2$. Consider input weights $(\mathbf{a}_{1}^{T},b_{1})=(1,0),(\mathbf{a}_{2}^{T},b_{2})=(0,1)$ and output weights $c_{11}=3,c_{12}=5,c_{21}=4,c_{22}=2$. In this example, for simplicity we took all output weights to be positive so that we only deal with positive zonotopes. The zonotopes of the two outputs will be two parallelograms with parallel edges, as illustrated in the Figure 6. The zonotope of the first output is generated by $c_{11}(\mathbf{a}_{1}^{T},b_{1})=3(1,0)=(3,0)$ and $c_{12}(\mathbf{a}_{2}^{T},b_{2})=5(0,1)=(0,5)$, and of the second output by $c_{21}(\mathbf{a}_{1}^{T},b_{1})=4(1,0)=(4,0)$ and $c_{22}(\mathbf{a}_{2}^{T},b_{2})=2(0,1)=(0,2)$. Say we want to reduce the hidden neurons to $K=1,\tilde{\mathbf{f}}=(\tilde{f_{1}})$. If we could approximate each output’s zonotope separately, we could simply apply the single output algorithm and approximate each parallelogram by its diagonal. However, these diagonals are not parallel to each other, and thus can not occur by a single hidden neuron $\tilde{f}_{1}$ with input weights $(\tilde{\mathbf{a}}_{1}^{T},\tilde{b}_{1})$. Instead, we have to choose a single common direction $(\tilde{\mathbf{a}}_{1}^{T},\tilde{b}_{1})$ for both output zonotopes. We can however choose a different magnitude for each output along this common direction. As will be presented in the algorithm below, for the common direction we choose the vector $(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i})=\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}=(0.5,0.5)$. For each output $j$, $\tilde{c}_{j1}$ is chosen so that the edge is as close to the diagonal as possible. Specifically, we choose $\tilde{c}_{11}=3+5=8$ and $\tilde{c}_{21}=2+4=6$. The approximation procedure can be seen in Figure 6(c).
@@
5.2.1 Non-iterative Algorithm for multi-output network
To tackle this problem of simultaneous zonotope approximation, Misiakos et al. (2022) proposed the ”Neural Path K-means” algorithm. Their approach to the problem is to use K-means clustering on the set of vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]\right\}$, and replace the vectors of each formed cluster by a single representative, the center of the cluster (i.e. the mean of the vectors of the cluster). Our approach also uses K-means to cluster the same vectors, but instead of taking the representative to be the center of the cluster, we form the representative using the following heuristic:
Suppose you have 2 neurons with input weights $(\mathbf{a}_{1}^{T},b_{1}),(\mathbf{a}_{2}^{T},b_{2})$ and output weights $\mathbf{C}_{:,1}^{T},\mathbf{C}_{:,2}^{T}$, that get clustered together. If $(\mathbf{a}_{1}^{T},b_{1})\approx(\mathbf{a}_{2}^{T},b_{2})\approx\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}$, then for every output $j$ the following holds:
$C_{j,1}(\mathbf{a}_{1}^{T},b_{1})+C_{j,2}(\mathbf{a}_{2}^{T},b_{2})\approx(C_{j,1}+C_{j,2})\frac{(\mathbf{a}_{1}^{T},b_{1})+(\mathbf{a}_{2}^{T},b_{2})}{2}.$
Thus, in order to approximate the 2 neurons, it suffices to take the mean of their input weights and the sum of their output weights. This heuristic can also be motivated by viewing at the network: it corresponds to taking the mean in terms of the input activations of the two neurons, and taking the sum as the joint output activations of the two neurons.
We generalize the above to clusters with more neurons. For every cluster $k\in[K]$ with clustered neuron indexes $I_{k}$ and vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\right\}$ take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of $\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\right\}$, take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of $\left\{\mathbf{C}_{:,i}^{T},i\in I_{k}\right\}$, and form the representative of the cluster $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$. This leads to Algorithm 2.
@@ 5.2.1 Non-iterative Algorithm for multi-output network
1:  Execute KMeans with $K$ centers on the vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})$ for $i\in[n]$, forming $K$ clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})\mid i\in I_{k}\}$ for $k\in[K]$.
@@
There are several variations of Algorithm 2. For instance, one could use the sum instead of the mean for the input weights. However, in our experiments, this approach yielded worse results. Another variation involves clustering neurons based on cosine similarity rather than the $L^{2}$ distance of the vectors, recognizing that neurons with parallel input weights can be clustered together regardless of their output weights. Additionally, one could normalize the vectors based on their input weights by shifting the norm to the output weights before applying K-means. Despite testing these methods, none provided significant advantages. Therefore, we simply recommend Algorithm 2.
5.2.2 Iterative Algorithm for multi-output network
To improve the approximation of Algorithm 2, we again make use of tropical geometry. Specifically, we formulate an optimization problem that takes the output of Algorithm 2 and with an iterative process produces weights that achieve a better simultaneous zonotope approximation.
Motivated by Algorithm 1, assuming the number of null neurons are few (see Misiakos et al. (2022) and Appendix A.6), we wish in terms of every output $j$ the cluster representative $\widetilde{C}_{j,k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be as close to the cluster sum $\sum_{i\in I_{k}}C_{j,i}(\mathbf{a}_{i}^{T},b_{i})$ as possible. Thus, for every cluster $k$, we have unknowns $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})=(\tilde{a}_{k1},\ldots,\tilde{a}_{kd},\tilde{b}_{k},\widetilde{C}_{1k},\ldots)$, and we wish to find a solution which minimizes the following criterion:
$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$
@@ 5.2.2 Iterative Algorithm for multi-output network
1. Holding the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ constant and minimizing with respect to $\widetilde{\mathbf{C}}_{:,k}$, we notice that the terms of the sum are independent and thus can be minimized separately. For each term of the sum, the minimization occurs if we project the sum $\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})$ onto $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$. We have:
$\widetilde{C}_{jk}=\frac{\left\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\rangle}{\left\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|^{2}}$
2. Holding the output weights $\widetilde{\mathbf{C}}_{:,k}$ constant, and minimizing with respect to the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$, we take the derivative of the criterion with respect to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and set it to zero. We have:
@@
$(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
We iterate over the above alternating steps for num_iter iterations. Additionally, we can initialize the iteration with the representative obtained from Algorithm 2, which should provide a good starting point. This initialization is expected to reduce the number of required epochs and speed up the compression algorithm. The resulting procedure is detailed in Algorithm 3. We should note that, in practise, the number of null generators is not negligible, and thus our optimization criterion also constitutes a heuristic method. For this reason, the number of iterations should not be excessive. We provide the following bound for the approximation error of our algorithm (proof in Appendix A.6).
@@
1:  Execute K-means with K centers on the vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]$, and form K clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\},k\in[K]$
2:  for $k\in[K]$ do
3:     Form the initial cluster representative $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ as follows:
4:      2i. Take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of the input weights + bias of the vectors of the cluster, i.e., the mean of the set of vectors $\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\}$
5:      2ii. Take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of the output weights of the vectors of the cluster, i.e., the sum of the set of vectors $\{\mathbf{C}_{:,i}^{T},i\in I_{k}\}$

### v3

@@ 3 Approximation based on Hausdorff distance
In this section, we present our refined theorem, which uses the Hausdorff distance in its standard continuous form to bound the error between two tropical polynomials. In the following, the distance between two points $\mathbf{u}$ and $\mathbf{v}$ is denoted as $\text{dist}(\mathbf{u},\mathbf{v}):=\|\mathbf{u}-\mathbf{v}\|$, where $\|\cdot\|$ denotes the standard $L^{2}$ Euclidean norm. The distance between a point $\mathbf{u}$ and a set $V$ is defined as $\operatorname{dist}(\mathbf{u},V)=\operatorname{dist}(V,\mathbf{u}):=\inf_{\mathbf{v}\in V}\|\mathbf{u}-\mathbf{v}\|$.
Definition 1 (Hausdorff distance).
Let $S,\tilde{S}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance H(S,S~)H($\tilde{S}$) of the two sets is defined as
In the case of polytopes $P,\tilde{P}$, due to their convexity and compactness, the suprema in the above expression are attained, and in fact by points in the vertex sets $V_{P},V_{\tilde{P}}$ of the polytopes.
@@
Theorem 2.
[…]
@@
Since $V_{P}\subseteq P,V_{\tilde{P}}\subseteq\tilde{P}$, we have that $\mathrm{dist}(\mathbf{u},\tilde{P})\leq\mathrm{dist}(\mathbf{u},V_{\tilde{P}}),\mathrm{dist}(P,\mathbf{v})\leq\mathrm{dist}(V_{P},\mathbf{v})$. The equality in the aforementioned inequalities is achieved only in special cases and thus the bound H(P,P~)H($\tilde{P}$) we provide is in general tighter than DH(P,P~)DH($\tilde{P}$) from Misiakos et al. (2022).
By applying the triangle inequality and Theorem 2 we obtain the following result for two neural networks of compatible input and output dimension.
Corollary 3.
Let $\mathbf{v}$ and $\tilde{\mathbf{v}}$ be the outputs of two neural networks as in Figure 1. Then, the following inequality holds:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right),$
@@ 4 Compression Algorithm
From Cor. 3 it is evident that to compress a ReLU network consisting of a pair of consecutive linear layers, one has to choose compressed weight matrices $\widetilde{\mathbf{A}}\in\mathbb{R}^{K\times(d+1)}$ and $\widetilde{\mathbf{C}}\in\mathbb{R}^{m\times K}$ with $K<n$ such that $\widetilde{P}_{j}\approx P_{j}$ and $\widetilde{Q}_{j}\approx Q_{j}$ for all $j\in[m]$, where the approximation relation is in terms of the Hausdorff distance between the zonotopes.
@@ 4.1 Single output
If we have $m=1$ then the problem can readily be translated into a zonotope approximation problem—the task of approximating a zonotope with another zonotope that has fewer generators. This problem is known as zonotope order reduction (Yang and Scott, 2018). In our case, the approximation must happen in terms of the Hausdorff distance.
Algorithm for single output network.
For single-output networks, our approach uses K-Means to cluster positive and negative generators, and replaces each cluster with a single representative. Unlike Misiakos et al. (2022), who define the representative as the center/mean of the cluster, we take the representative to be the sum of the cluster.
Our Algorithm 1 is depicted below. Figure 2 illustrates an example execution of TropNNC compared to the method of Misiakos et al. (2022). For our algorithm, the following bound and its corollary hold.
@@
$\{$ $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}):i\in C_{+}\},$
$\{$ $|\tilde{c}_{i}|(\tilde{\mathbf{a}}_{i}^{T},\tilde{b}_{i}):i\in C_{-}\},$
@@
5: For each constructed neuron $i$ set the output weight to 11 if the neuron corresponds to a representative of a positive cluster ($i\in C_{+}$), otherwise set it to -1.
(a) Original
(b) Misiakos et al.
@@
Figure 2: Example execution of TropNNC compared to the method of Misiakos et al. (2022).
Proposition 4.
@@
[…]
@@ 4.2 Multi-output
We now consider the multi-output case with $m\in\mathbb{N}$. Notice an interesting property of the zonotopes $P_{j},Q_{j}$: they share the directions $(\mathbf{a}_{i}^{T},b_{i})$ of their generators. For instance, output $v_{1}$ might have a positive generator $|c_{1i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $P_{1}$, while output $v_{2}$ might have a negative generator $|c_{2i}|(\mathbf{a}_{i}^{T},b_{i})$ of zonotope $Q_{2}$. These generators are parallel to each other, with common direction $(\mathbf{a}_{i}^{T},b_{i})$. Hence, our original positive and negative zonotopes have parallel generators, and we aim to approximate them with new positive and negative zonotopes, also with parallel generators. We refer to this complex approximation problem as simultaneous zonotope approximation.
Example 1.
[…]
@@
4.2.1 Non-iterative Algorithm for multi-output network.
For the multi-output case, we perform clustering of similar neurons based on the set of clustering vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in[n]\right\}$, and replace each cluster with a single representative. Unlike Misiakos et al. (2022), who take the cluster center/mean as the representative, we form the representative as follows: For every cluster $k\in[K]$ with clustered neuron indexes $I_{k}$ and vectors $\left\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T}),i\in I_{k}\right\}$ take $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be the mean of $\left\{(\mathbf{a}_{i}^{T},b_{i}),i\in I_{k}\right\}$, take $\widetilde{\mathbf{C}}_{:,k}^{T}$ to be the sum of $\left\{\mathbf{C}_{:,i}^{T},i\in I_{k}\right\}$, and form the representative of the cluster $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$. The complete procedure is shown in Algorithm 2.
@@
1: Execute clustering on the clustering vectors $(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})$ for $i\in[n]$, forming $K$ clusters $\{(\mathbf{a}_{i}^{T},b_{i},\mathbf{C}_{:,i}^{T})\mid i\in I_{k}\}$ for $k\in[K]$.
@@
4.2.2 Iterative Algorithm for multi-output network.
To improve the approximation of Algorithm 2, we formulate an optimization problem that takes the output of Algorithm 2 and with an iterative process produces weights that achieve a better simultaneous zonotope approximation.
Motivated by Algorithm 1, assuming the number of null neurons are few (see appendix), we want in terms of every output $j$ the cluster representative $\widetilde{C}_{j,k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ to be a close approximation to the cluster sum $\sum_{i\in I_{k}}C_{j,i}(\mathbf{a}_{i}^{T},b_{i})$. Thus, for every cluster $k$, we have unknowns $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})=(\tilde{a}_{k1},\ldots,\tilde{a}_{kd},\tilde{b}_{k},\widetilde{C}_{1k},\ldots)$, and we wish to find a solution which minimizes the following criterion:
$\sum_{j=1}^{m}\left\|\tilde{C}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})-\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\right\|^{2},$ (1)
@@
1. Fixing the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and minimizing with respect to $\widetilde{\mathbf{C}}_{:,k}$, the terms of the sum are independent and thus can be minimized separately. For each term of the sum, the optimal error occurs if we project the sum $\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})$ onto $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$. We have:
$\widetilde{C}_{jk}=\frac{\left\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\rangle}{\left\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|^{2}}$ (2)
2. Fixing the output weights $\widetilde{\mathbf{C}}_{:,k}$, and minimizing with respect to the input weights $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$, we take the derivative of criterion (1) with respect to $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ and set it to zero. We have:
@@
$(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})=\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$ (3)
We can initialize the iteration with the representative obtained from Algorithm 2. The resulting procedure is detailed in Algorithm 3. We note that, in practice, the number of null generators is not negligible, and thus criterion (1) also constitutes a heuristic method. For this reason, the number of iterations should not be excessive. We provide Proposition 6 for the approximation error of our algorithm.
@@
1: Form the initial cluster representatives $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k},\widetilde{\mathbf{C}}_{:,k}^{T})$ by executing the non-iterative algorithm.
2:for $\text{iter}=1\text{ to }num\_iter$ do
3:  for $k\in[K]$ do
4:   for $j\in[m]$ do
5:    $\widetilde{C}_{jk}\leftarrow\frac{\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\rangle}{\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\|^{2}}$